# On Championing Foundation Models: From Explainability to Interpretability

### v1

@@ 3.1.1 Interpreting Through Generalization Analysis
In this line of research examining gradient flow behavior over the loss function, the primary focus remains on linear problem settings. To rigorously analyze gradient flow dynamics, studies like [110] often simplify the transformer architecture to a single-layer attention mechanism with linear activation functions. Specifically, [110] assumes that the parameters of a linear task are drawn from a Gaussian prior, $N(0,\psi^{2}\mathbf{I}_{d})$, and the prompt $C_{n}={(x_{1},y_{1}),\dots,(x_{n},y_{n})}$ is an independent copy of the testing sample $(x,y)$, where $x\sim N(0,\mathbf{H})$, $\mathbf{H}\succeq 0$, and $y\sim N(\beta^{T}x,\sigma^{2})$. Furthermore, they consider a pretraining dataset where each of the $T$ independent regression tasks is associated with $n+1$ in-context prompt examples. The generalization is defined as $\mathcal{R}_{fixed}(f_{M})=\mathbb{E}[(y-f_{M}(C_{n},x;\theta))^{2}]$, establishing a sample complexity bound for the pretraining of a linear parameterized single-layer attention model within the framework of linear regression under a Gaussian prior. The theorem is stated as follows:
@@
(Task complexity for pertaining [110]). Let $n\geq 0$ represent the quantity of in-context examples utilized to generate the dataset. Let $\theta_{T}$ denote the parameters output by SGD on the pretraining dataset with a geometrically decaying step size. Assume that the initialization $\theta_{0}$ commutes with $\mathbf{H}$ and $\gamma_{0}\leq 1/(c\operatorname{tr}(\mathbf{H})\operatorname{tr}(\tilde{\mathbf{H}}_{n}))$, where $c>1$ is an absolute constant and $\tilde{\mathbf{H}}_{n}$ is defined as $\psi^{2}\mathbf{H}((\mathrm{tr}(\mathbf{H})/n+\sigma^{2}/n\psi^{2})\mathbf{I}+(n+1)/n\mathbf{H})$. Then we have
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\lesssim\left\langle\mathbf{H}\tilde{\mathbf{H}}_{n},\left(\prod_{t=1}^{T}\left(\mathbf{I}-\gamma_{t}\mathbf{H}\tilde{\mathbf{H}}_{n}\right)\theta_{n}^{*}\right)^{2}\right\rangle+\left(\psi^{2}\mathbf{tr}(\mathbf{H})+\sigma^{2}\right)\frac{D_{\text{eff}}}{T_{\text{eff}}}.$ (2)
@@
$T_{\text{eff }}:=\frac{T}{\log(T)},\quad D_{\text{eff }}:=\sum_{i}\sum_{j}\min\left\{1,T_{\text{eff }}^{2}\gamma_{0}^{2}\lambda_{i}^{2}\tilde{\lambda}_{j}^{2}\right\}$
@@
$\tilde{\lambda}_{j}:=\psi^{2}\lambda_{j}\left(\frac{\operatorname{tr}(\mathbf{H})+\sigma^{2}/\psi^{2}}{n}+\frac{n+1}{n}\lambda_{j}\right),\quad j\geq 1$
@@
Another significant line of research employs statistical learning theory techniques, including covering numbers, PAC-Bayes arguments, and algorithm stability [105, 109, 108, 93]. Specifically, [105] and [109] utilize chaining arguments with covering numbers to derive generalization bounds. In [105], the authors employ standard uniform concentration analysis via covering numbers to establish an excess risk guarantee. Specifically, they consider embedding dimensions $D$, the number of layers $L$, the number of attention heads $M$, feedforward width $D^{\prime}$, and the norm of parameters $\|\theta\|\leq B$. Furthermore, they explore the use of $T$ prompts where each has $|C_{n+1}|=n+1$ in-context examples for pretraining, leading to the derivation of the following excess risk bound:
@@
$\mathcal{R}_{fixed}(f_{M}(C_{n},x;\hat{\theta}))-\operatorname*{min}_{\theta}\mathcal{R}_{fixed}(f_{M}(C_{n},x;\theta))\lesssim\sqrt{\frac{L^{2}\left(MD^{2}+DD^{\prime}\right)\log B}{T}}.$
@@
$\mathcal{R}_{fixed}(f_{M}(C_{n},x;\hat{\theta}))\lesssim\sigma^{2}+\left(\sqrt{\frac{d^{4}}{T}}+\frac{s\log d}{n}\right)$
It is important to highlight that the theoretical result in Theorem 3.3 of research [105] deviates from the standard transformer architecture by substituting the softmax activation in the attention layers with a normalized ReLU function. This differs from the findings in Theorem 3.2 of research [133], which adheres to the standard architecture. Furthermore, [109] also employs the standard transformer architecture with softmax activation, focusing on a more specific context involving sequential decision-making problems. It utilizes covering arguments to derive a generalization upper bound of $\mathcal{O}(1/\sqrt{n})$ for the average regret.
Additionally, under PAC-Bayes framework, [108] investigate the pretraining error and draw a link between the pretraining error and ICL regret. Their findings show that the total variation distance between the trained model and a reference model is bounded as follows:
@@
$\begin{split}&\boldsymbol{X}_{(i)}=\text{ParallelMLPs}\left(\operatorname{ATTN}\left(\boldsymbol{X}_{(i-1)}\right)\right)\ \text{where}\ \operatorname{ATTN}(\boldsymbol{X}):=\operatorname{softmax}\left(\boldsymbol{X}\boldsymbol{W}_{i}\boldsymbol{X}^{\top}\right)\boldsymbol{X}\boldsymbol{V}_{i}\end{split}$
@@
Additionally, we need to provide further remarks regarding this theorem [93]. While the reliance on depth is growing exponentially, this constraint is less significant for most transformer architectures, as they typically do not have extreme depths. For instance, from 12 to 48 layers are present in various versions of GPT-2 and BERT [134, 135]. In this theorem, the upper bound on $\Gamma$ ensures that no single token can exert significant influence over another. The crucial technical aspect of this result is demonstrating the stability of the self-attention layer, which is the core component of a transformer. This stability is essential for maintaining the integrity and performance of the model, as it ensures that the influence of any individual token is limited, thereby preserving the overall structure and coherence of the model’s output.
Assume the model is trained on $T$ tasks, each including a data sequence with $n$ examples. Each sequence is input to the model in an auto-regressive manner during training. By conceptualizing ICL as an algorithmic learning problem, [93] derive a multitask generalization rate of $1/\sqrt{nT}$ for i.i.d. data. To establish the appropriate reliance on sequence length (the $\sqrt{n}$ factor), they address temporal dependencies by correlating generalization with algorithmic stability. Next, we introduce the excess risk bound, which quantifies the performance gap between the learned model and the optimal model. This measure accounts for both generalization error and optimization error, providing a comprehensive evaluation of the model’s effectiveness.
@@
For all $T$ tasks, let’s assume that the set of algorithmic hypotheses $\mathcal{A}$ is $\frac{K}{n}$-stable as defined in 2.4, and that the loss function $\ell(y,\cdot)$ is $L$-Lipschitz continuous, taking values in the interval $[0,1]$. Allow $\mathcal{A}$ to stand for the empirical solution that was acquired. Afterwards, the following bound is satisfied by the excess multi-task learning (MTL) test risk with a probability of at least $1-2\delta$:

### v2

@@ 3.4 Dynamic Behavior Analysis
Next, let us recall that our objective is to minimize the risk $\mathcal{R}(f)$ through optimization techniques. However, since the data distribution $\mathcal{D}$ is unknown, we cannot directly optimize this risk function. Instead, we approximate it by minimizing the empirical risk $\mathcal{R}_{n}(f)$ using methods such as gradient descent and stochastic gradient descent (SGD).
Gradient descent is an iterative optimization method where, at each step, we refine the solution by moving in the direction opposite to the gradient of the function at the current point. In contrast, SGD addresses this limitation by allowing the optimization process to proceed along a randomly selected direction, provided the expected direction aligns with the negative gradient.
Below, we will first introduce the fundamentals of the gradient descent algorithm and discuss the associated convergence analysis. The gradient of a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ at a point $\boldsymbol{\theta}$, represented as $\nabla f(\boldsymbol{\theta})$, is a vector consisting of $f$’s partial derivatives with respect to each component of $\boldsymbol{\theta}$:
$\nabla f(\boldsymbol{\theta})=\left(\frac{\partial f(\boldsymbol{\theta})}{\partial\theta_{1}},\ldots,\frac{\partial f(\boldsymbol{\theta})}{\partial\theta_{d}}\right).$
Gradient descent proceeds iteratively, beginning from an initial point $\boldsymbol{\theta}$, often set to zero (e.g., $\boldsymbol{\theta}^{(1)}=0$). In each iteration, the algorithm updates the current point by moving in the direction opposite to the gradient. Specifically, the update rule is:
$\boldsymbol{\theta}^{(t+1)}=\boldsymbol{\theta}^{(t)}-\eta\nabla f(\boldsymbol{\theta}^{(t)}),$
where $\eta>0$ is the step size. Intuitively, because the gradient indicates the direction of the steepest ascent for $f$ at $\boldsymbol{\theta}^{(t)}$, taking a small step in the opposite direction reduces $f$’s value. After $T$ iterations, the algorithm produces an output that approximates the minimum. We explore the convergence rate of gradient descent by using a convex and Lipschitz continuous function as an example.
Theorem 3.10.
(Convergence rate of GD for convex-Lipschitz functions [66]) Consider a convex function $f$ that is $\rho$-Lipschitz, and let $\boldsymbol{\theta}^{\star}\in\arg\min_{\{\boldsymbol{\theta}:\|\boldsymbol{\theta}\|\leq B\}}f(\boldsymbol{\theta})$ denote an optimal solution within the constraint $\|\boldsymbol{\theta}\|\leq B$. When we apply gradient descent to $f$ for $T$ iterations with a step size $\eta=\sqrt{\frac{B^{2}}{\rho^{2}T}}$, the resulting average vector $\bar{\boldsymbol{\theta}}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\theta^{(t)}}$ satisfies the inequality
$f(\bar{\boldsymbol{\theta}})-f(\boldsymbol{\theta}^{\star})\leq\frac{B\rho}{\sqrt{T}}.$
Moreover, to ensure that $f(\bar{\boldsymbol{\theta}})-f(\boldsymbol{\theta}^{\star})\leq\epsilon$ for any $\epsilon>0$, it is sufficient to perform gradient descent for a number of iterations satisfying
$T\geq\frac{B^{2}\rho^{2}}{\epsilon^{2}}.$
Proof Sketch for Convergence Rate of Gradient Descent for Convex-Lipschitz Functions in Theorem 3.10
This proof framework establishes an upper bound on the convergence rate of gradient descent when applied to a convex function $f$ that is also $\rho$-Lipschitz. Given a constraint on the norm of the optimal solution $\|\boldsymbol{\theta}^{\star}\|\leq B$, the objective is to show that after $T$ iterations with an appropriately chosen step size $\eta$, the average iterate achieves a suboptimality gap that decreases at the rate $O\left(\frac{1}{\sqrt{T}}\right)$.
Step 1: Definition of Suboptimality and Jensen’s Inequality Application.
  • Suboptimality Definition: We aim to bound the suboptimality of the averaged iterate $\bar{\boldsymbol{\theta}}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\theta}^{(t)}$ with respect to the optimal solution $\boldsymbol{\theta}^{\star}$,i.e.,
    $f(\bar{\boldsymbol{\theta}})-f(\boldsymbol{\theta}^{\star}).$
  • Jensen’s Inequality: Since $f$ is convex, we can apply Jensen’s inequality to show:
    $f(\bar{\boldsymbol{\theta}})-f(\boldsymbol{\theta}^{\star})=f\left(\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\theta}^{(t)}\right)-f(\boldsymbol{\theta}^{\star})\leq\frac{1}{T}\sum_{t=1}^{T}\left(f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star})\right).$
    This reduces the problem to bounding the average of the suboptimalities $f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star})$ over the iterations.
Step 2: Utilizing Convexity to Relate Suboptimality and Gradients .
  • Convexity Inequality: By convexity of $f$, for each $t$,
    $f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star})\leq\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\nabla f(\boldsymbol{\theta}^{(t)})\rangle.$
    This allows us to bound the suboptimality at each iteration in terms of the inner product of the difference $\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star}$ with the gradient $\nabla f(\boldsymbol{\theta}^{(t)})$.
  • Reduction to Summing Gradient Inner Products: Combining the above results, we find that
    $f(\bar{\boldsymbol{\theta}})-f(\boldsymbol{\theta}^{\star})\leq\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\nabla f(\boldsymbol{\theta}^{(t)})\rangle.$
    Define $\boldsymbol{v}_{t}=\nabla f(\boldsymbol{\theta}^{(t)})$ for convenience.
Step 3: Expanding Inner Products via the Update Rule.
  • Inner Product Expansion: Observe that
    $\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle=\frac{1}{\eta}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\eta\boldsymbol{v}_{t}\rangle$
    $=\frac{1}{2\eta}\left(-\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star}-\eta\boldsymbol{v}_{t}\|^{2}+\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star}\|^{2}+\eta^{2}\|\boldsymbol{v}_{t}\|^{2}\right).$
    Using the definition of the update rule $\boldsymbol{\theta}^{(t+1)}=\boldsymbol{\theta}^{(t)}-\eta\boldsymbol{v}_{t}$, we can rewrite this as:
    $\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle=\frac{1}{2\eta}\left(-\|\boldsymbol{\theta}^{(t+1)}-\boldsymbol{\theta}^{\star}\|^{2}+\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star}\|^{2}\right)+\frac{\eta}{2}\|\boldsymbol{v}_{t}\|^{2}.$
Step 4: Summing Over Iterations and Telescoping
  • Telescoping Sum: Summing this inequality over $t=1,\ldots,T$, we get:
    $\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle$
    $=\frac{1}{2\eta}\sum_{t=1}^{T}\left(-\|\boldsymbol{\theta}^{(t+1)}-\boldsymbol{\theta}^{\star}\|^{2}+\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star}\|^{2}\right)+\frac{\eta}{2}\sum_{t=1}^{T}\|\boldsymbol{v}_{t}\|^{2}.$ (2)
    The first term on the right side is a telescoping sum that simplifies to:
    $\|\boldsymbol{\theta}^{(1)}-\boldsymbol{\theta}^{\star}\|^{2}-\|\boldsymbol{\theta}^{(T+1)}-\boldsymbol{\theta}^{\star}\|^{2}.$
  • Simplified Bound: Substituting back, we get:
    $\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\leq\frac{1}{2\eta}\|\boldsymbol{\theta}^{(1)}-\boldsymbol{\theta}^{\star}\|^{2}+\frac{\eta}{2}\sum_{t=1}^{T}\|\boldsymbol{v}_{t}\|^{2}.$
    Assuming $\boldsymbol{\theta}^{(1)}=0$, we further simplify to:
    $\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\leq\frac{1}{2\eta}\|\boldsymbol{\theta}^{\star}\|^{2}+\frac{\eta}{2}\sum_{t=1}^{T}\|\boldsymbol{v}_{t}\|^{2}.$
Step 5: Choosing Step Size and Final Convergence Rate
  • Step Size Selection: Set $\eta=\sqrt{\frac{B^{2}}{\rho^{2}T}}$ so that each gradient $\|\boldsymbol{v}_{t}\|\leq\rho$ results in
    $\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\leq\frac{B\rho}{\sqrt{T}}.$
  • Final Bound: This choice of step size ensures that after $T$ iterations,
    $f(\bar{\boldsymbol{\theta}})-f(\boldsymbol{\theta}^{\star})\leq\frac{B\rho}{\sqrt{T}}.$
    To achieve an error $\epsilon$, it is sufficient to perform $T\geq\frac{B^{2}\rho^{2}}{\epsilon^{2}}$ iterations.
In SGD, the update direction does not need to precisely align with the true gradient. Instead, we permit the update direction to be a random vector, requiring only that its expected value aligns with the gradient direction at each iteration. More generally, this expectation must be consistent with the subgradient of the function at the current point in the optimization process. The update for $\boldsymbol{\theta}^{(t+1)}$ is:
$\boldsymbol{\theta}^{(t+1)}=\boldsymbol{\theta}^{(t)}-\eta\boldsymbol{v}_{t},$
where $\eta$ is the step size, and $\boldsymbol{v}_{t}$ is chosen randomly from a distribution such that $\mathbb{E}[\boldsymbol{v}_{t}\mid\boldsymbol{\theta}^{(t)}]\in\partial f(\boldsymbol{\theta}^{(t)})$, ensuring that, on average, the update direction remains within the subdifferential of $f$ at $\boldsymbol{\theta}^{(t)}$. Given that the expected value of $\boldsymbol{v}_{t}$ corresponds to the subgradient of $f$ at $\mathbf{w}^{(t)}$, it remains possible to establish a bound on the expected outcome of stochastic gradient descent, similar to that in Theorem 3.10. This result is rigorously formalized in the following theorem.
Theorem 3.11.
(Convergence rate of SGD for convex-Lipschitz functions [66]) Let $B,\rho>0$, and let $f$ be a convex and $\rho$-Lipschitz function with an optimal solution $\boldsymbol{\theta}^{\star}\in\arg\min_{\|\boldsymbol{\theta}\|\leq B}f(\boldsymbol{\theta})$. Suppose that we run SGD for $T$ iterations, using a step size $\eta=\sqrt{\frac{B^{2}}{\rho^{2}T}}$, and assume that $\|\boldsymbol{v}_{t}\|\leq\rho$ holds with probability 1 for each iteration $t$. Then we have
$\mathbb{E}[f(\bar{\boldsymbol{\theta}})]-f(\boldsymbol{\theta}^{\star})\leq\frac{B\rho}{\sqrt{T}}.$
Thus, to ensure $\mathbb{E}[f(\bar{\boldsymbol{\theta}})]-f(\boldsymbol{\theta}^{\star})\leq\epsilon$ for any $\epsilon>0$, it is sufficient to run the SGD algorithm for a number of iterations satisfying
$T\geq\frac{B^{2}\rho^{2}}{\epsilon^{2}}.$
Proof Sketch for Convergence Rate of SGD for Convex-Lipschitz Functions in Theorem 3.11.
Step 1: Taking Expectation of the Suboptimality Bound
  • Expectation of Suboptimality: The goal is to bound the expected suboptimality of the averaged iterate $\bar{\boldsymbol{\theta}}=\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{\theta}^{(t)}$ with respect to the optimal solution $\boldsymbol{\theta}^{\star}$, i.e.,
    $\mathbb{E}[f(\bar{\boldsymbol{\theta}})]-f(\boldsymbol{\theta}^{\star}).$
  • Applying Proof framework of Theorem 3.10: For any sequence $\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{T}$, we can obtain a bound on the inner product of the gradient with the difference $\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star}$. By taking expectations, we find:
    $\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\right]\leq\frac{B\rho}{\sqrt{T}}.$
Step 2: Establishing a Comparison Between Expected Suboptimality and Gradient Inner Products.
  • Expected Suboptimality Inequality: It remains to show that
    $\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\frac{1}{T}\sum_{t=1}^{T}\left(f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star})\right)\right]\leq\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\right].$
  • Using Linearity of Expectation: By the linearity of expectation, we rewrite the right-hand side as:
    $\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\frac{1}{T}\sum_{t=1}^{T}\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\right]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\right].$
Step 3: Applying the Law of Total Expectation.
  • Law of Total Expectation: For two random variables $\alpha$ and $\beta$, and a function $g$, the law of total expectation states $\mathbb{E}_{\alpha}[g(\alpha)]=\mathbb{E}_{\beta}\mathbb{E}_{\alpha}[g(\alpha)|\beta]$. Applying this with $\alpha=\boldsymbol{v}_{1:t}$ and $\beta=\boldsymbol{v}_{1:t-1}$, we find:
    $\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\right]=\mathbb{E}_{\boldsymbol{v}_{1:t-1}}\left\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\mathbb{E}_{\boldsymbol{v}_{t}}\left[\boldsymbol{v}_{t}\mid\boldsymbol{v}_{1:t-1}\right]\right\rangle.$
    Since $\boldsymbol{\theta}^{(t)}$ depends only on $\boldsymbol{v}_{1:t-1}$, once we condition on $\boldsymbol{v}_{1:t-1}$, the vector $\boldsymbol{\theta}^{(t)}$ becomes deterministic.
Step 4: Applying the SGD Condition on Expected Gradients.
  • SGD Condition on Gradients: The property of SGD ensures that $\mathbb{E}_{\boldsymbol{v}_{t}}[\boldsymbol{v}_{t}\mid\boldsymbol{\theta}^{(t)}]\in\partial f(\boldsymbol{\theta}^{(t)})$, meaning $\mathbb{E}_{\boldsymbol{v}_{t}}[\boldsymbol{v}_{t}\mid\boldsymbol{v}_{1:t-1}]\in\partial f(\boldsymbol{\theta}^{(t)})$.
  • Convexity and Expected Suboptimality: Since $f$ is convex, we have:
    $\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\mathbb{E}_{\boldsymbol{v}_{t}}\left[\boldsymbol{v}_{t}\mid\boldsymbol{v}_{1:t-1}\right]\rangle\geq f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star}).$
    Taking the expectation with respect to $\boldsymbol{v}_{1:t-1}$, we conclude that:
    $\mathbb{E}_{\boldsymbol{v}_{1:t}}\left[\langle\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\star},\boldsymbol{v}_{t}\rangle\right]\geq\mathbb{E}_{\boldsymbol{v}_{1:t-1}}\left[f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star})\right].$
Step 5: Concluding the Proof with Expected Suboptimality Bound.
  • Summing Over All Iterations: Summing over $t$, dividing by $T$, and applying the linearity of expectation yields:
    $\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[f(\boldsymbol{\theta}^{(t)})-f(\boldsymbol{\theta}^{\star})\right]\leq\mathbb{E}_{\boldsymbol{v}_{1:T}}\left[\frac{B\rho}{\sqrt{T}}\right].$
The proof framework is complete.
@@ 4.1.1 Interpreting Through Generalization Analysis
In this line of research examining gradient flow behavior over the loss function, the primary focus remains on linear problem settings. To rigorously analyze gradient flow dynamics, studies like [90] often simplify the transformer architecture to a single-layer attention mechanism with linear activation functions. Specifically, [90] assumes that the parameters of a linear task are drawn from a Gaussian prior, $N(0,\psi^{2}\mathbf{I}_{d})$, and the prompt $C_{n}={(x_{1},y_{1}),\dots,(x_{n},y_{n})}$ is an independent copy of the testing sample $(x,y)$, where $x\sim N(0,\mathbf{H})$, $\mathbf{H}\succeq 0$, and $y\sim N(\beta^{T}x,\sigma^{2})$. Furthermore, they consider a pretraining dataset where each of the $T$ independent regression tasks is associated with $n+1$ in-context prompt examples. The generalization is defined as
$\mathcal{R}_{fixed}(f_{M})=\mathbb{E}[(y-f_{M}(C_{n},x;\theta))^{2}],$
establishing a sample complexity bound for the pretraining of a linear parameterized single-layer attention model within the framework of linear regression under a Gaussian prior. The theorem is stated as follows:
@@
(Task complexity for pertaining [90]). Let $n\geq 0$ represent the quantity of in-context examples utilized to generate the dataset. Let $\theta_{T}$ denote the parameters output by SGD on the pretraining dataset with a geometrically decaying step size. Assume that the initialization $\theta_{0}$ commutes with $\mathbf{H}$ and $\gamma_{0}\leq 1/(c\operatorname{tr}(\mathbf{H})\operatorname{tr}(\tilde{\mathbf{H}}_{n}))$, where $c>1$ is an absolute constant and $\tilde{\mathbf{H}}_{n}$ is defined as
$\psi^{2}\mathbf{H}((\mathrm{tr}(\mathbf{H})/n+\sigma^{2}/n\psi^{2})\mathbf{I}+(n+1)/n\mathbf{H}).$
Then we have
@@
$\lesssim\left\langle\mathbf{H}\tilde{\mathbf{H}}_{n},\left(\prod_{t=1}^{T}\left(\mathbf{I}-\gamma_{t}\mathbf{H}\tilde{\mathbf{H}}_{n}\right)\theta_{n}^{*}\right)^{2}\right\rangle+\left(\psi^{2}\mathbf{tr}(\mathbf{H})+\sigma^{2}\right)\frac{D_{\text{eff}}}{T_{\text{eff}}}.$ (3)
@@
$T_{\text{eff }}:=\frac{T}{\log(T)},\quad D_{\text{eff }}:=\sum_{i}\sum_{j}\min\left\{1,T_{\text{eff }}^{2}\gamma_{0}^{2}\lambda_{i}^{2}\tilde{\lambda}_{j}^{2}\right\},$
@@
$\tilde{\lambda}_{j}:=\psi^{2}\lambda_{j}\left(\frac{\operatorname{tr}(\mathbf{H})+\sigma^{2}/\psi^{2}}{n}+\frac{n+1}{n}\lambda_{j}\right),\quad j\geq 1.$
@@
Another significant line of research employs statistical learning theory techniques, including covering numbers, PAC-Bayes arguments, and algorithm stability [91, 85, 87, 86]. Specifically, [91] and [85] utilize chaining arguments with covering numbers to derive generalization bounds. In [91], the authors employ standard uniform concentration analysis via covering numbers to establish an excess risk guarantee. Specifically, their technique adopts the chaining-based proof framework detailed in Theorem 3.5 of Section 3.1. They investigate key architectural parameters, including embedding dimensions $D$, the number of layers $L$, the number of attention heads $M$, feedforward width $D^{\prime}$, and the norm of parameters $\|\theta\|\leq B$. Furthermore, they explore the use of $T$ prompts where each has $|C_{n+1}|=n+1$ in-context examples for pretraining, leading to the derivation of the following excess risk bound:
@@
$\mathcal{R}_{fixed}(f_{M}(C_{n},x;\hat{\theta}))-\operatorname*{min}_{\theta}\mathcal{R}_{fixed}(f_{M}(C_{n},x;\theta))$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\lesssim\sqrt{\frac{L^{2}\left(MD^{2}+DD^{\prime}\right)\log B}{T}}.$
@@
$\mathcal{R}_{fixed}(f_{M}(C_{n},x;\hat{\theta}))\lesssim\sigma^{2}+\left(\sqrt{\frac{d^{4}}{T}}+\frac{s\log d}{n}\right).$
It is important to highlight that the theoretical result in Theorem 4.3 of research [91] deviates from the standard transformer architecture by substituting the softmax activation in the attention layers with a normalized ReLU function. This differs from the findings in Theorem 4.2 of research [90], which adheres to the standard architecture. Furthermore, [85] also employs the standard transformer architecture with softmax activation, focusing on a more specific context involving sequential decision-making problems. It utilizes covering arguments, as demonstrated within the proof framework of Theorem 3.5, to derive a generalization upper bound of $\mathcal{O}(1/\sqrt{n})$ for the average regret.
Additionally, [87] investigate the pretraining error and establish a connection between the pretraining error and the ICL regret. Their technique primarily builds upon the PAC-Bayes proof framework we outlined for Theorem 3.9. Their findings demonstrate that the total variation distance between the trained model and a reference model is bounded as follows:
@@
$\boldsymbol{X}_{(i)}=\text{ParallelMLPs}\left(\operatorname{ATTN}\left(\boldsymbol{X}_{(i-1)}\right)\right),$
where
$\operatorname{ATTN}(\boldsymbol{X}):=\operatorname{softmax}\left(\boldsymbol{X}\boldsymbol{W}_{i}\boldsymbol{X}^{\top}\right)\boldsymbol{X}\boldsymbol{V}_{i}.$
@@
Additionally, we need to provide further remarks regarding this theorem [86]. While the reliance on depth is growing exponentially, this constraint is less significant for most transformer architectures, as they typically do not have extreme depths. For instance, from 12 to 48 layers are present in various versions of GPT-2 and BERT [94]. In this theorem, the upper bound on $\Gamma$ ensures that no single token can exert significant influence over another. The crucial technical aspect of this result is demonstrating the stability of the self-attention layer, which is the core component of a transformer. This stability is essential for maintaining the integrity and performance of the model, as it ensures that the influence of any individual token is limited, thereby preserving the overall structure and coherence of the model’s output.
Assume the model is trained on $T$ tasks, each including a data sequence with $n$ examples. Each sequence is input to the model in an auto-regressive manner during training. By conceptualizing ICL as an algorithmic learning problem, [86] derive a multitask generalization rate of $1/\sqrt{nT}$ for i.i.d. data. To establish the appropriate reliance on sequence length (the $\sqrt{n}$ factor), they address temporal dependencies by correlating generalization with algorithmic stability. Their approach synthesizes the stability bound proof framework, as rigorously detailed in Theorem 3.8, with the covering number analysis presented in Theorem 3.5. By combining these complementary methodologies, they establish a solid theoretical foundation that addresses both stability and capacity, enabling a detailed analysis of the model’s generalization behavior. Next, we introduce the excess risk bound, which quantifies the performance gap between the learned model and the optimal model. This measure accounts for both generalization error and optimization error, providing a comprehensive evaluation of the model’s effectiveness.
@@
For all $T$ tasks, let’s assume that the set of algorithmic hypotheses $\mathcal{A}$ is $\frac{K}{n}$-stable as defined in 3.7, and that the loss function $\ell(y,\cdot)$ is $L$-Lipschitz continuous, taking values in the interval $[0,1]$. Allow $\mathcal{A}$ to stand for the empirical solution that was acquired. Afterwards, the following bound is satisfied by the excess multi-task learning (MTL) test risk with a probability of at least $1-2\delta$: